OpenStudy (anonymous):

Please check!!!!! (8x+1)^1/2 =5 ANS:3

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

The answer choices are ::: 3 2 -3 none of the above

OpenStudy (anonymous):

You can check your answer easily by setting 3 as x

OpenStudy (anonymous):

When you do that, you get 5 on the left side

OpenStudy (anonymous):

oh my god. im too tired. i shouldve known hahaha -____- thanks haha

OpenStudy (anonymous):

No worries. Merry Christmas

OpenStudy (anonymous):

What anime is that in your picture? I need an anime to watch...

OpenStudy (anonymous):

haha it's from tumblr. I can find out if you'd like!

OpenStudy (anonymous):

@NewOne

OpenStudy (anonymous):

could you help with another?

OpenStudy (anonymous):

and this is not a no brainer.. haha

OpenStudy (anonymous):

Sure. I'm only playing wii right now

OpenStudy (anonymous):

lol oh .. sorry to interrupt

OpenStudy (anonymous):

\[3\sqrt{x-3}-2=16\]

OpenStudy (anonymous):

x = 6 x = 9 x = 39 None of the above

OpenStudy (anonymous):

Here is a tip for these questions. Since the right side is pretty big, you want the result of the square root pretty big as well.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

39?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

\[\sqrt{x+3}-3=x \] would this be -2 and -3?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

fan freakin tastic. thanks i'll hit you up, for further help!

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

\[(x ^{1/2})^{1/6}\] what about this?

OpenStudy (anonymous):

Sorry to be a bother

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

how would you simplify that?

OpenStudy (anonymous):

x^1/4??

OpenStudy (anonymous):

Nope. \[(x^{a})^{b} = x^{ab} \]

OpenStudy (anonymous):

x^1/12 then?

OpenStudy (anonymous):

you multiply denominators?

OpenStudy (anonymous):

You can refer the chart here for exponential laws. http://www.mathsisfun.com/algebra/exponent-laws.html I find this page helpful whenever I forget them.

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

thank you!